A flashlight has 6 batteries, 2 of which are defective. If 2 are selected at random without replacement, find the probability that both are defective.
Answer by Alan3354(69443)

The 1st is 2 of 6.
The 2nd is 1 of 5.
= 1/3*1/5 = 1/15